The records of Norton, Inc. show the following for July.

Standard labor-hours allowed per unit of output

2.5

Standard variable overhead rate per standard direct labor-hour

$40

Good units produced

60,000

Actual direct labor-hours worked

151,000

Actual total direct labor

$7,649,000

Direct labor efficiency variance

$50,000 U

Actual variable overhead

$5,839,000

Required:

Compute the direct labor and variable overhead price and efficiency variances. (Do not round intermediate calculations. Indicate the effect of each variance by selecting “F” for favorable, or “U” for unfavorable. If there is no effect, do not select either option.)

Direct labor:

   Price variance

   Efficiency variance

Variable overhead:

   Price variance

   Efficiency variance
Answer #1

Labor efficiency variance = (SH-AH) *SR

(2.5*60,000 - 151,000)*SR = - 50,000

(150,000 - 151,000)*SR = - 50,000

SR = 50

Labor price Variance = (SP-AP) *AH

= (50 - 7,649,000/151,000)*151,000

= 99,000 U

Labor efficiency variance

= 50,000 U

Variable overhead price Variance = (SP-AP) *AH

= (40 - 5,839,000/151000)*151,000

= 201,000 F

Variable overhead efficiency variance = (SH-AH) *SR

= (60,000*2.5 - 151,000)*40

= 40,000 U